40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 28 20:42:53 2020, Koji, Update, General, Vent Prep 10x
    Reply  Thu Oct 29 12:43:35 2020, gautam, Update, General, Vent Prep preVentAlignment.pngPRMIcarr_PRMalignment.pngALS_ool.pdflowPowerMC.png
       Reply  Thu Oct 29 19:30:56 2020, gautam, Update, General, Vent complete ventValveStatus.pngvent_Oct2020.jpgventValveStatus_overnight.pnglowPowerIMC_afterVent.png
          Reply  Mon Nov 2 12:44:25 2020, gautam, Update, General, ETMY and OMC chamber heavy doors off CB84FD62-C48B-414F-8825-50C9DE0B47CF.jpeg
             Reply  Mon Nov 2 16:46:06 2020, gautam, Update, General, Plan for OMC chamber IMG_2318.JPGIMG_2332.JPG
                Reply  Mon Nov 2 17:13:19 2020, Koji, Update, General, Plan for OMC chamber 
                   Reply  Mon Nov 2 17:32:05 2020, gautam, Update, General, Plan for OMC chamber IMG_2317.JPG
Message ID: 15653     Entry time: Mon Nov 2 12:44:25 2020     In reply to: 15652     Reply to this: 15654
Author: gautam 
Type: Update 
Category: General 
Subject: ETMY and OMC chamber heavy doors off 

[koji, rana, gautam]

This morning, we did the following;

  1. Removed the ETMY chamber heavy door. It is stored on the rack at the east end.
  2. Removed OMC chamber heavy door. It is smaller than the other chamber doors, so doesn't sit on the standard size rack we have. So it is laid flat, on a clean sheet of ameristat, on a cart next to the NE corner of the PSL table.
  3. After taking some photos of the chamber and making sure the position of the suspension tower was marked by some stops, we removed the ETMY cage and moved it to the cleanroom area. The optic was EQ stopped during the transport, and the OSEMs were removed. The ETMY Oplev HeNe was turned off and the PSL was shuttered to allow us to work without goggles.
  4. The broken off magnet was retrieved from inside the OSEM. The shadow sensor voltage recovered a value of ~800cts, which means the LED/PD pair on the UR OSEM seems to work fine, and it was in fact the magnet blocking the PD that was the issue.
  5. In the cleanroom area, we removed the optic from the wire loop and placed it on the magnet gluing fixture. The wire is intact (for now), so there is some hope of re-suspending it in the same loop.

The OSEMs remain in the EY vacuum chamber. The next set of steps are:

  1. Clean the EP30-2 residue from the broken magnet joint - this will require some scrubbing with an acetone soaked scrub or similar implement.
  2. Reglue the broken magnet.

We will most likely work on this tomorrow. At ~1615, I briefly opened the PSL shutter and tweaked the IMC alignment. We will almost certainly change the pointing into the IMC when we remove the old OMC and rebalance that table, so care should be taken when working on that...

Quote:
  1. We are now ready to take the doors off. 
Attachment 1: CB84FD62-C48B-414F-8825-50C9DE0B47CF.jpeg  212 kB  Uploaded Mon Nov 2 14:12:27 2020  | Hide | Hide all
CB84FD62-C48B-414F-8825-50C9DE0B47CF.jpeg
ELOG V3.1.3-